Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Yi Qiang
Log entry time 08:23:52 on December 03, 2008
Entry number 251836
This entry is a followup to: 251673
keyword=Experiment Summary for Dec 2, Owl - Swing
Summary for Dec 1 (owl - swing)
Smooth data taking.
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/12/01 owl | 6.74 |
0.10 | 1.17 |
2008/12/01 day | 0.00 |
0.00 | 8.00 |
2008/12/01 swing | 0.00 |
0.00 | 8.00 |
total | 6.74 |
0.10 | 17.17 |
Main events/issues:
- Production before 7:30am.
- Beam study/RF cavity recovery/Beam recorvery afterwards.
- Put in new beam line window collimator for BigBite during down time.
Plans for the next few days:
- Production.
- Change HRS polarity during Friday Day shift.